OpenStudy (anonymous):

what exactly is a linear equation and when would you use them in a word problem? Examples please.

OpenStudy (anonymous):

Linear equations are those in which the dependent variable is directly proportional to 1st power of the independent variable. Also you should not see terms involving squares of the independent variable. and also the dependent variable is not a function of a function of independent variable. eg of 1st order equations : 1. V = IR (not I has power = 1) 2. S = V/T equations which are not 1st order : 1. V=I^2 2. V = I*R*f(I) (where f(I) is function of I)

OpenStudy (anonymous):

alright, but what is a function?

OpenStudy (anonymous):

if you don't know, it does not matter for your question. If you want to know anyway, I suggest you post a separate question.

OpenStudy (anonymous):

a function is a equation in simple terms. like in V =IR V is said to be a function of I. It's like a way to say how out put is related to input. V is out put and I is in put. R in this case is constant. (just for additional info, V=IR is ohms law and R is an experimentally determined constant)

OpenStudy (anonymous):

A function is not just an equation......